# Text-Guided Multi-Property Molecular Optimization with a Diffusion Language Model

### v1

@@ 5.2. Metrics
Following the previous work outlined by Edwards et al.  (Edwards et al., 2022), we employed eight metrics to assess the structural construction abilities and four criteria to evaluate property optimization capabilities.
@@
  • MACCS FTS  (Durant et al., 2002), RDK FTS  (Schneider et al., 2015), and Morgan FTS  (Rogers and Hahn, 2010). These metrics calculate the average Tanimoto similarity between the fingerprints of the optimized and target molecules.
@@
  • FCD  (Preuer et al., 2018). Fréchet ChemNet Distance (FCD) evaluates the latent information agreement.
  • ADMET properties accuracy. According to the dataset, the property changes are categorized into high-¿low, low-¿high, and no change. A property prediction model (Jiang et al., 2023) trained on ADMET properties in the MMP dataset is utilized to predict the optimized molecules’ corresponding attributes. Then, the accuracy of property change type between the optimized and source molecules is calculated to convey the model’s optimization ability toward the desired orientation.
@@ 5.3. Baselines
    Chemformer  (Irwin et al., 2022). This is a transformer-based model pre-trained on unlabeled SMILES and finetuned for MO task.
@@
It is worth noting that MIMOSA, Modof, MolSearch, and FRATTVAE optimize molecules guided by the external property predictors.
@@
5.4. Experiment Settings
The SMILES vocabulary included 265 tokens, with token embeddings being trainable at $d=32$. We used a pre-trained SciBERT model  (Beltagy et al., 2019) with an embedding dimension of $d_{1}=768$ to encode the textual descriptions. Moreover, the core transformer architecture, $f_{\theta}$, consists of $L=12$ layers with a hidden dimension of $d_{2}=1024$. Our TransDLM has around 181 million trainable parameters.
For the MO task, we used a uniform step-skipping strategy during the denoising process, which reduced the sampling steps from 1,000 to 200. As a result, optimizing a molecule takes approximately 1.02 seconds on an AMD EPYC 7763 (64 cores) @ 2.450GHz CPU and a single NVIDIA A6000 GPU. During training, we set the total diffusion steps to $T=2000$. The Adam optimizer’s  (Kingma, 2014) learning rate was $5e^{-5}$, and we included a linear warm-up.
5.5. Overall Performance
@@
As shown in Tables  1 and  2, TransDLM generally outperformed all baseline models. Although some baselines excelled in validity, they performed poorly in other metrics, which demonstrates limitations. Specifically, TransDLM achieved the best BLEU score and an 8.6% improvement in Levenshtein distance, indicating better alignment with the target molecules’ structures and higher sequence accuracy. Most advantages on FTS criteria further highlighted its ability to capture molecular fingerprints, providing deeper insights into the molecule’s structural information. Notably, we observed a 43.8% improvement in FCD, which indicates that TransDLM generated molecules with closer distributions to the training data, thereby outperforming other methods. Though Chemformer matched more targets in Exact score, it failed to perform well in property requirements. Regarding ADMET properties, drastic LogD, Solubility, and Clint improvements by 30.8%, 1.8% and 0.3%, resulted in a remarkable 27.9% increase in the ratio of optimized molecules that met all ADMET property criteria.
@@ 5.6. Ablation Study
We performed two ablation studies to validate the effectiveness of the TransDLM strategies, as outlined below:
@@
As shown in Tables  1 and  2, optimization guided by descriptions made of IUPAC names displayed a better performance than of original SMILES strings. This implies that IUPAC nomenclature conveys more physical and chemical semantics information which benefits text-guided MO. Similarly, though sampling from the encoded source molecules sacrificed few advantages on property criteria, it generated superior outcomes for most structural metrics than from pure noise, indicating that our sampling strategy optimizes molecules without destroying the original scaffolds.
@@ 5.7. Case Study Analysis
In the overall performance, the TransDLM model demonstrated robust physical construction and chemical optimization capabilities. To validate that out approach not only exhibited improvements in numerical performance but also genuinely satisfies the MO goal of moderate structural disparities and substantial chemical dissimilarities, we conducted a case study analysis as illustrated in Fig.  4.
In the first two cases, the two MMPs underwent one atom and one functional group substitution. TransDLM precisely identified the site that needed replacement and carried out structural modifications, fulfilling all ADMET property requirements under the guidance of the textual descriptions. In cases where TransDLM did not fully recognize all modification sites, such as the third case, TransDLM also maintained the core scaffold of the source molecule effectively, fulfilling the requirements of chemical properties without overly modifying the physical structure. In contrast, other methods either significantly disrupted the original molecule’s structure or failed to meet the chemical properties required by MMPs, underscoring their limitations.
@@ 6. CONCLUSION
In this paper, we proposed TransDLM, a novel language diffusion model for text-guided SMILES-based MO, which utilizes IUPAC nomenclature for molecular semantic representation and optimizes molecules guided by physically and chemically detailed textual descriptions. Instead of relying on an external property predictor, error propagation caused by the predictor was prevented. Additionally, we theoretically demonstrated TransDLM’s effectiveness and performed validation experiments using the MMP dataset. Notably, TransDLM outperformed other state-of-the-art methods on most structural metrics and all ADMET property criteria. Furthermore, TransDLM enables simultaneous sampling of multiple molecules, making it highly suitable for scalable web-based applications and allows for efficient large-scale optimization, leveraging the web’s capacity for distributed computation. Furthermore, these achievements occurred without relying on additional data sources or pre-training, underscoring the effectiveness of TransDLM in text-guided multi-property MO.

### v3

@@ 5.2. Metrics
Following the previous work outlined by Edwards et al. (Edwards et al., 2022), we employed eight metrics to assess the structural construction abilities and six criteria to evaluate property optimization capabilities.
@@
  • MACCS (Durant et al., 2002), RDK (Schneider et al., 2015), and Morgan FTSs (Rogers and Hahn, 2010). These metrics calculate the average Tanimoto similarity between the fingerprints of the optimized and target molecules.
@@
  • FCD (Preuer et al., 2018) metric. The FCD metric leverages the penultimate layer of a pre-trained network called ChemNet, which incorporates chemical and biological information for more sophisticated comparisons.
  • Accuracy of ADMET property changes. According to the dataset, the property changes are categorized into decrease, increase, and remain. A property prediction model (Jiang et al., 2023) trained on ADMET properties in the MMP dataset is utilized to predict the optimized molecules’ corresponding properties. Then, the accuracy of property change type between the optimized and source molecules is calculated to convey the model’s optimization ability toward the desired orientation.
  • HV and R2 indicator. The hypervolume indicator (HV) (Zitzler et al., 2003) measures the volume covered by the Pareto front solutions in the objective space relative to a reference point, while the R2 indicator (Brockhoff et al., 2012) assesses the Pareto front’s performance under different preference weightings. We use these indicators to determine whether the optimized molecules are a desirable solution set on quantitative estimate of drug-likeness (QED) and synthetic accessibility score (SA).
@@ 5.3. Baselines
    DyMol (Shin et al., [n. d.]). This is a divide-and-conquer approach combined with a decomposition strategy for multi-property optimization.
It is worth noting that all competing approaches optimize molecules guided by the external property predictors.
@@
5.4. Overall performance and visualization
@@
As shown in Tables 1 and 2, TransDLM mostly outperformed all baseline models across multiple evaluation metrics, demonstrating its superiority in multi-property MO. Specifically, TransDLM achieved the best BLEU score and an 11.8 % improvement in Levenshtein distance, indicating better alignment with the target molecules’ structures and higher sequence accuracy. Most advantages on FTS criteria further highlighted its ability to capture molecular fingerprints, providing deeper insights into the molecule’s structural information. Notably, we observed a 78 % improvement in FCD, which indicates that TransDLM generated molecules with closer distributions to the training data, thereby outperforming other methods. Regarding ADMET properties, drastic LogD, Solubility, and Clint improvements by 57 %, 2.4 % and 0.3 %, resulted in a remarkable 46.7 % increase in the ratio of optimized molecules that met all ADMET property criteria. Additionally, TransDLM outperformed all competing methods across HV and R2 indicators on QED and SA, indicating our approach can stably and comprehensively generate high-quality and practically feasible candidate molecules. These results highlight TransDLM’s advantage in balancing structural fidelity with functional optimization, making it a more effective approach for multi-property MO. Notably, detailed statistical indicator testing is described in Appendix C.
For the Validity metric, several baselines obtain a perfect score because they rely on iterative search over discrete molecular graphs or fragments and explicitly discard invalid candidates. This guarantees validity at the cost of higher computational overhead. In contrast, TransDLM generates molecules in a single pass within a continuous latent space, making it more time-efficient. Our model achieves a validity of 0.993, and the few invalid cases stem from the stochastic nature of continuous generation. Although extended training can slightly improve validity, we found it compromises other critical metrics such as property optimization and structural similarity, resulting in an unfavorable trade-off.
To further validate that TransDLM genuinely satisfies the MO goal of moderate structural disparities (Zhou et al., 2019; He et al., 2022) and substantial chemical dissimilarities, we visualized some representative results in Fig. 4.
In the first two examples, the two MMPs underwent one atom and one functional group substitution. TransDLM precisely identified the site that needed replacement and carried out structural modifications, fulfilling all ADMET property requirements under the guidance of the textual descriptions. In examples where TransDLM did not fully recognize all modification sites, such as the third example, TransDLM also maintained the core scaffold of the source molecule effectively, fulfilling the requirements of chemical properties without overly modifying the physical structure. In contrast, other methods either significantly disrupted the original molecule’s structure or failed to meet the chemical properties required by MMPs, underscoring their limitations.
Table 3. Optimization results for structural similarity compared with ablation studies on the test split of MMP dataset. Additionally, $\mathrm{TransDLM}_{\mathrm{noise}}$ and $\mathrm{TransDLM}_{\mathrm{SMILES}}$ respectively denote sampling from pure noise and optimizing guided by SMILES string-based descriptions.
[…]
TransDLM 0.74 0.009 14.838 0.818 0.792 0.665 0.109 0.993
BERT 0.723 0.004 15.771 0.725 0.709 0.611 0.134 0.897
BioBERT 0.737 0.012 13.92 0.796 0.754 0.673 0.115 0.997
SciBERT 0.74 0.009 14.838 0.818 0.792 0.665 0.109 0.993
Table 4. Optimization results for ADMET properties compared with ablation studies on the test split of MMP dataset.
[…]
$\mathrm{TransDLM}_{\mathrm{noise}}$ 0.16 0.789 0.757 0.108 0.927 0.124
$\mathrm{TransDLM}_{\mathrm{SMILES}}$ 0.038 0.756 0.753 0.025 0.912 0.127
TransDLM 0.157 0.783 0.757 0.11 0.929 0.121
BERT 0.122 0.682 0.752 0.083 0.92 0.139
BioBERT 0.125 0.779 0.755 0.104 0.921 0.137
SciBERT 0.157 0.783 0.757 0.11 0.929 0.121
Figure 5. Visualization of XAC and optimized XAC binding to A1R and A2AR. The lower the docking score, the greater the binding affinity.
@@ 5.5. Ablation study
We performed three ablation studies to validate the effectiveness of the TransDLM strategies. as outlined below:
@@
  • Different pre-trained language models in TransDLM were selected for ablation studies: BERT (Devlin et al., 2019), BioBERT (Lee et al., 2020), and SciBERT (Beltagy et al., 2019).
As shown in Tables 3 and 4, optimization guided by descriptions made of IUPAC names displayed a better performance than of original SMILES strings. This implies that IUPAC nomenclature conveys more physical and chemical semantics information which benefits text-guided MO. Similarly, though sampling from the encoded source molecules sacrificed few advantages on property criteria, it generated superior outcomes for most structural metrics than from pure noise, indicating that our sampling strategy optimizes molecules without destroying the original scaffolds.
The ablation study conducted on multiple pre-trained language models revealed that BERT consistently underperformed across all metrics, while BioBERT and SciBERT demonstrated comparable on structural similarity metrics, but SciBERT yields the best property optimization results.
This performance gap primarily arises from the differences in their pre-training corpora. BERT is pre-trained on general-domain texts, such as Wikipedia and various books, which limits its exposure to scientific terminology, particularly IUPAC nomenclature and property descriptions. As a result, BERT’s ability to accurately interpret chemical semantics is constrained. In contrast, BioBERT leverages biomedical literature and thus better captures drug-related concepts. However, its focus on biology and pharmacology results in weaker coverage of fundamental chemical principles.
In comparison, SciBERT is pre-trained on a broader scientific corpus that encompasses computer science and an abundance of fundamental chemical principles along with experimental descriptions. Consequently, SciBERT exhibits a deeper understanding of essential chemical concepts such as IUPAC nomenclature and chemical reactions when compared to BioBERT.
@@ 5.6. Case study analysis
In this section, we present the application of TransDLM through a case study involving XAC, a ligand recognized for its dual binding affinity to A1R and A2AR receptors (Cheng et al., 2017). The primary objective of this optimization was to modulate the binding affinities of XAC to enhance its selectivity toward the A1R receptor while reducing its affinity for A2AR. The optimization process entailed optimizing XAC conforming to the specified binding affinity textual description. To validate the effectiveness of the optimization, we utilized Schrödinger [^4] https://www.schrodinger.com, a professional molecular docking software, to evaluate the docking scores between both the original XAC molecule and its optimized counterpart with the A1R and A2AR receptors.
The yielded docking scores aligned well with our expectations, as illustrated in Fig. 5. Specifically, the optimized XAC showed a significantly higher docking score for A1R and a lower one for A2AR compared to the original molecule. This outcome confirms that our optimization successfully achieved the desired preferential binding. Detailed experimental scheme is available in Appendix D.
Notably, optimizing ligand-receptor binding affinity is an out-of-distribution (OOD) task for TransDLM, which was trained on the ADMET-related dataset. Its success demonstrates that our text-guided framework possesses strong generalization capabilities and flexibility. We can reasonably infer that if TransDLM were to be trained on a broader and more diverse dataset which includes binding affinity data, it has the potential to evolve into a powerful optimization tool applicable across various drug discovery scenarios, extending beyond mere ADMET properties.
6. Limitations and future work
Although TransDLM achieves promising results, several limitations remain and point to directions for future research.
A key limitation lies in its multi-stage pipeline. The current framework involves RECAP decomposition, IUPAC parsing, and pre-trained language model embedding before the diffusion process. This modular design improves controllability and interpretability but also increases the system complexity compared to end-to-end models. The reliance on external tools, especially for IUPAC conversion, adds computational overhead and complicates large-scale deployment. Future work will aim to streamline these steps through deeper integration or end-to-end trainable components capable of generating structured guidance directly from molecular inputs.
Another challenge is the computational burden of diffusion-based generation. Like other diffusion models, TransDLM requires iterative denoising, which is more expensive than single-step generative methods like Generative Adversarial Networks (GANs) or VAEs. While this trade-off yields superior diversity and quality, efficiency remains a concern. We have already reduced sampling steps from 2000 to 200 using a step-skipping strategy with minimal performance loss. Future acceleration could be achieved through techniques like knowledge distillation (Hinton et al., 2015) or Denoising Diffusion GANs (Xiao et al., 2022), enabling more scalable deployment in industrial settings.
Additionally, we plan to extend TransDLM to broader property optimization tasks, including non-biological properties such as QED and SA, as well as affinity toward specific protein targets. We also aim to improve generalizability by training on larger and more diverse molecular libraries, enhancing robustness on OOD tasks. These efforts will further strengthen TransDLM’s applicability in drug discovery and chemical engineering domains, paving the way for more efficient and targeted molecular optimization.
@@ 7. Conclusion
In this paper, we proposed TransDLM, a novel diffusion language model for text-guided MO, which utilizes IUPAC nomenclature for molecular semantic representation and optimizes molecules guided by physically and chemically detailed textual descriptions. To retain the core scaffolds of source molecules, TransDLM samples from the encoded source molecules. Moreover, instead of relying on an external property predictor, error propagation was greatly mitigated during diffusion process. Additionally, we theoretically demonstrated TransDLM’s effectiveness and performed validation experiments. Notably, TransDLM outperformed other methods on most structural metrics and all ADMET property criteria, demonstrating robust structural construction and property enhancement capabilities. Notably, the successful optimization of XAC using TransDLM underscores its effectiveness and potential in out-of-distribution problems. Furthermore, these achievements occurred without relying on additional data sources or pre-training, underscoring the effectiveness of TransDLM in text-guided multi-property molecular optimization.
Code and data availability
The code and data used in this study are now publicly available at: https://github.com/Cello2195/TransDLM.
Acknowledgement
This work was supported in part by the Natural Science Foundation of China (No.62476203), and the Guangdong Provincial Natural Science Foundation General Project (No.2025A1515012155). Engineering Research Center for Big Data Application in Private Health Medicine of Fujian Universities, Putian University, Putian, Fujian351100, China (MKF202405). Key Program of Hubei Natural Science Foundation Traditional Chinese Medicine Innovation and Development Joint Fund (2025AFD470).
Appendix A Details of experimental settings
Our TransDLM has around 181 million trainable parameters, implemented with PyTorch 2.0.1 (Paszke et al., 2019) and CUDA 11.7. As a result, optimizing a molecule takes approximately 1.02 seconds on an AMD EPYC 7763 (64 cores) @ 2.450GHz CPU and a single NVIDIA A6000 GPU. Table 5 demonstrates all the hyper-parameters of TransDLM.
Appendix B Details of RECAP disassembly
RECAP is a molecular disassembly function built in RDKit. Specifically, RECAP emulates the forward synthesis process conducted in a laboratory to perform reverse operations, executing a series of conversions and decompositions on molecules, ultimately yielding a set of plausible molecular fragments. RECAP is capable of tracking the disassembly process and constructing a tree-like data structure. The original molecule is designated as the root node, while the disassembled molecule serves as the parent node. Molecules obtained from disassembly are recorded as child nodes; all branch nodes under a specific node are classified as descendant nodes, and all parent nodes associated with a particular node are referred to as ancestor nodes. Finally, molecules that cannot be further disassembled are identified as leaf nodes.
Appendix C Statistical Indicator Testing
To assess whether the performance improvements of our method over competing baselines are statistically significant rather than due to randomness, we conducted statistical indicator testing based on repeated experiments.
Figure 6. Statistical indicator testing across baselines on 5 critical metrics. The dashed lines indicate p-value thresholds of 0.05 and 0.001.
We chose 5 critical metrics for statistical indicator testing, BLEU, MACCS, FCD, ALL and HV, which comprehensively cover syntactic similarity, substructure properties, chemical space similarity , multi-property optimization success and Pareto front quality. For each method, we performed 5 independent runs using different random seeds under identical experimental settings. Additionally, paired t-test served as the primary statistical test and Benjamini-Hochberg (FDR-BH) (Benjamini and Hochberg, 1995) was used to correct for the false discovery rate across the set of comparisons. As illustrated in Fig. 6, the results of our analysis are compelling: out of the 25 corrected p-values (5 metrics $\times$ 5 baselines), 24 were less than 0.05, and 22 were less than 0.001.
Appendix D Details of case study analysis
Referring to the experimental configuration of existing research (Li et al., 2025b), we located potential binding sites utilizing Schrödinger’s SiteMap tool (Halgren, 2007) with a cubic exploration grid ($15\AA \times 15\AA \times 15\AA$ ) centered on the receptor’s critical functional areas. The XAC was processed through LigPrep, where protonation states were generated at pH 7.0 $\pm$ 2.0 using Epik (up to 32 states per ligand) (Shelley et al., 2007). Molecular docking was subsequently carried out with the Glide module in Maestro (Friesner et al., 2004), and docking scores for each pose were automatically computed to identify the most favorable binding conformations.